# Accelerated Convergence of Frank–Wolfe Algorithms with Adaptive Bregman Step-Size Strategy

### v1

@@ 4     $x_{t+1}\leftarrow(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$
[…]
any weak quad. [^3] We assume that $f$ is local quadratic growth. $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(\epsilon^{-1/\nu})$ (Thm. 5.5)
[…]
any [^5] Its rate is established for the Frank–Wolfe gap. nonconvex $\mathcal{O}(\epsilon^{-2})$ [Lacoste-Julien2016-mj] $\mathcal{O}(\epsilon^{-1-1/\nu})$ (Thm. 5.1)
Outline
The structure of this paper is as follows. Section 2 introduces essential notations, including Bregman distances, the smooth adaptable property, weak convexity, growth conditions, and scaling inequalities, which are crucial for the convergence analyses that come later. In Section 3, we present the FW algorithm with an adaptive Bregman step-size strategy and the away-step FW algorithm utilizing the Bregman distance. Section 4 details the convergence results for convex optimization, demonstrating both sublinear and linear convergence rates. Section 5 addresses convergence results for the nonconvex case, showing that our algorithm achieves global convergence to a stationary point of (1.1). Under the assumption that $f$ is weakly convex and local quadratic growth, we establish local linear convergence to a minimizer of (1.1), indicating that the proposed algorithms converge to a minimizer when the initial point is sufficiently close. Section 6 presents applications to nonnegative linear inverse problems, phase retrieval, low-rank minimization, and nonnegative matrix factorization (NMF).
2 Preliminaries
[…]
Let $\operatorname{int}{C}$, $\operatorname{ri}{C}$, and $\operatorname{cl}{C}$ be the interior, the relative interior, and the closure of a set $C\subset\mathbb{R}^{n}$, respectively. The set $\operatorname{conv}C$ denotes the convex hull of $C$, and the set $\operatorname{Vert}C$ denotes the set of vertices of $C$. We also define the distance from a point $x\in\mathbb{R}^{n}$ to $C$ by $\operatorname{dist}(x,C):=\inf_{y\in C}\|x-y\|$.
For an extended-real-valued function $f:\mathbb{R}^{n}\to[-\infty,+\infty]$, the effective domain of $f$ is defined by $\operatorname{dom}f:=\{x\in\mathbb{R}^{n}\,|\,f(x)<+\infty\}$. The function $f$ is said to be proper if $f(x)>-\infty$ for all $x\in\mathbb{R}^{n}$ and $\operatorname{dom}f\neq\emptyset$. Let $[f\leq\zeta]:=\{x\in\mathbb{R}^{n}\,|\,f(x)\leq\zeta\}$ be a $\zeta$-sublevel set of $f$ for some $\zeta\in\mathbb{R}$. Let $\mathcal{C}^{k}$ be the class of $k$-times continuously differentiable functions for $k\geq 0$. The sign function $\operatorname{sgn}(x)$ is defined by $\operatorname{sgn}(x)=-1$ for $x<0$, $\operatorname{sgn}(x)=0$ for $x=0$, and $\operatorname{sgn}(x)=1$ for $x>0$.
2.1 Bregman Distances
Let $C$ be a nonempty open convex subset of $\mathbb{R}^{n}$.
  1. (i)
    $\phi$ is proper, lower semicontinuous, and convex, with $\operatorname{dom}\phi\subset\operatorname{cl}C$ and $\operatorname{dom}\partial\phi=C$.
  2. (ii)
Definition 2.2 (Bregman Distance [Bregman1967-rf]).
Given a kernel generating distance $\phi\in\mathcal{G}(C)$, a Bregman distance $D_{\phi}:\operatorname{dom}\phi\times\operatorname{int}\operatorname{dom}\phi\to\mathbb{R}_{+}$ associated with $\phi$ is defined by
$D_{\phi}(x,y)=\phi(x)-\phi(y)-\langle\nabla\phi(y),x-y\rangle.$ (2.1)

### v3

@@ 3.1 Bregman Step-Size Strategy for $\gamma_{t}$
2      Choose $\beta<1$, $\eta\leq 1$, and $\tau>1$
3      $M\leftarrow\eta\tilde{L}$, $\kappa\leftarrow 1$
4      while $D_{f}(x_{+},x)\leq M\gamma^{1+\kappa}D_{\phi}(v,x)$ do
5           $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x),x-v\rangle}{M(1+\kappa)D_{\phi}(v,x)}\right)^{1/\kappa},\gamma_{\max}\right\}$
6           $x_{+}\leftarrow(1-\gamma)x+\gamma v$
7           $M\leftarrow\tau M$
8           $\kappa\leftarrow\beta\kappa$
9
10           $\tilde{L}^{*}\leftarrow M$, $\tilde{\nu}^{*}\leftarrow\kappa$
11           return $\tilde{L}^{*}$, $\tilde{\nu}^{*}$, $\gamma$
12
Algorithm 2 Adaptive Bregman step-size strategy
If $\phi=\frac{1}{2}\|\cdot\|^{2}$, we have $\nu=1$ and $\gamma_{t}=\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|v_{t}-x_{t}\|^{2}}$, which is often called the (Euclidean) short step-size. Although the short step-size does not require line searches, it requires knowledge of the value of $L$.
Adaptive Bregman Step-Size:
The exact value or the tight upper bound of $L$ is often unknown, however, the algorithm’s performance heavily depends on it; an underestimation of $L$ might lead to non-convergence, and an overestimation of $L$ might lead to slow convergence. Moreover, a worst-case $L$ might be too conservative for regimes where the function is better behaved. Due to all these reasons, Pedregosa et al. [Pedregosa2020-ay] proposed an adaptive step-size strategy for FW algorithms in the case where $\phi=\frac{1}{2}\|\cdot\|^{2}$. We present our algorithm in Algorithm 2. This step-size strategy can be used as a drop-in replacement. For example, inserting $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},1)$ between lines 1 and 1 in Algorithm 1, we obtain that Algorithm 2 searches $L$ and $\nu$ satisfying (2.2).
Remark 3.5 (Well-definedness and termination of Algorithm 2).
By the $L$-smad property of $(f,\phi)$ and Lemma 2.4, we know that (2.2) holds for all $M\geq L$ and $\nu\geq\kappa>0$. Therefore, the loop in line 2 in Algorithm 2 is well-defined and guaranteed to terminate. It suffices to update $\kappa$ in line 2 only if (2.1) does not hold. Thus, $\kappa$ cannot become too small.
3.2 Bregman Away-Step Frank–Wolfe Algorithm
Input: $x_{0}\in\operatorname*{argmin}_{v\in P}\langle\nabla f(x),v\rangle$ for $x\in P$, $\beta<1$
1 $\mathcal{S}_{0}\leftarrow\{x_{0}\}$, $\lambda_{x_{0},0}\leftarrow 1$
2
3for $t=0,\ldots$ do
4      $v_{t}^{\textup{FW}}\leftarrow\operatorname*{argmin}_{v\in P}\langle\nabla f(x_{t}),v\rangle$, $v_{t}^{\textup{A}}\leftarrow\operatorname*{argmax}_{v\in\mathcal{S}_{t}}\langle\nabla f(x_{t}),v\rangle$
5
8
9           else
10                $v_{t}\leftarrow v_{t}^{\textup{A}}$, $d_{t}\leftarrow v_{t}^{\textup{A}}-x_{t}$, $\gamma_{t,\max}\leftarrow\lambda_{v_{t}^{\textup{A}},t}/(1-\lambda_{v_{t}^{\textup{A}},t})$
11
12                $\gamma_{t}\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},\gamma_{t,\max}\right\}$, $x_{t+1}\leftarrow x_{t}-\gamma_{t}d_{t}$
13
16                     $\lambda_{v_{t}^{\textup{FW}},t+1}\leftarrow\begin{cases}\gamma_{t}&\text{if}\ v_{t}^{\textup{FW}}\notin\mathcal{S}_{t}\\ (1-\gamma_{t})\lambda_{v_{t}^{\textup{FW}},t}+\gamma_{t}&\text{if}\ v_{t}^{\textup{FW}}\in\mathcal{S}_{t}\end{cases}$, $\mathcal{S}_{t+1}\leftarrow\begin{cases}\mathcal{S}_{t}\cup\{v_{t}^{\textup{FW}}\}&\text{if}\ \gamma_{t}<1\\ \{v_{t}^{\textup{FW}}\}&\text{if}\ \gamma_{t}=1\end{cases}$
17
18                     else
19                          $\lambda_{v,t+1}\leftarrow(1+\gamma_{t})\lambda_{v,t}$ for all $v\in\mathcal{S}_{t}\setminus\{v_{t}^{\textup{A}}\}$, $\lambda_{v_{t}^{\textup{A}},t+1}\leftarrow(1+\gamma_{t})\lambda_{v_{t}^{\textup{A}},t}-\gamma_{t}$
21
22
In the case where $P$ is a polytope, the classical FW algorithm might zigzag when approaching the optimal face and, in consequence, converge slowly. Away-step FW algorithm overcomes this drawback. Away steps allow the algorithm to move away from vertices in a convex combination of $x_{t}$, which can effectively shortcircuit the zigzagging. The convergence properties of this algorithm were unknown for a long time, and it was only quite recently that Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] established the linear convergence of the away-step FW algorithm. Inspired by [GueLat1986-bg, Lacoste-Julien2015-gb, Wolfe1970-ax], we propose a variant of the away-step FW algorithm utilizing Bregman distances as given in Algorithm 3.
For the following discussion, we introduce some notions. In the same way [Braun2022-ew], we define an active set $\mathcal{S}\subset\operatorname{Vert}P$, where $\operatorname{Vert}P$ denotes the set of vertices of $P$, and the away vertex as $v_{t}^{\textup{A}}\in\operatorname*{argmax}_{v\in\mathcal{S}}\langle\nabla f(x_{t}),v\rangle$, where $x_{t}$ is a (strict) convex combination of elements in $\mathcal{S}$, i.e., $x_{t}=\sum_{v\in\mathcal{S}}\lambda_{v}x_{t}$ with $\lambda_{v}>0$ for all $v\in\mathcal{S}$ and $\sum_{v\in\mathcal{S}}\lambda_{v}=1$. If $\nu$ is unknown, we can add line search procedures to search $\nu$ until (2.1) holds. If $\nu$ and $L$ are unknown, we can use $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$ with $\gamma_{t}\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{M(1+\kappa)D_{\phi}(v_{t},x_{t})}\right)^{1/\kappa},\gamma_{t,\max}\right\}$ instead of line 2 in Algorithm 2. When $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu\equiv 1$, Algorithm 3 corresponds to the Euclidean away-step FW algorithm.
4 Linear Convergence for Convex Optimization
In this section, we assume that $f$ is convex.
Assumption 4.1.
The objective function $f$ is convex.
We will now establish faster convergence rates than $O(1/t)$ up to linear convergence depending on the choice of parameters. First, we establish the faster convergence of Algorithm 1 with the Bregman short step-size, i.e., $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$, in the case where the optimal solution lies in the relative interior. Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$.
Theorem 4.2 (Linear convergence of FW algorithm with short step-size or adaptive step-size).
Suppose that Assumptions 3.1 and 4.1 hold. Let $\nu>0$ and let $f$ satisfy the HEB condition (2.3) with $q\geq 1+\nu$ and $\mu>0$ and $D$ be the diameter of $P$. Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$. Consider the iterates of Algorithm 1 with short step-size $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$ or the adaptive Bregman step-size strategy (Algorithm 2). Then, it holds that, for all $t\geq 1$,
$f(x_{t})-f^{*}\leq\begin{cases}\max\left\{\frac{1}{1+\nu},1-\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{c^{1+1/\nu}D^{2/\nu}}\right\}^{t-1}LD^{2}&\text{if}\ q=1+\nu,\\ \frac{LD^{2}}{(1+\nu)^{t-1}}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)(c/r)^{1+\nu}D^{2})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}(t-t_{0})\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
where $c=(q/\mu)^{1/q}$ and $t_{0}:=\max\{\lfloor\log_{1/(1+\nu)}((L(1+\nu(c/r)^{1+\nu}D^{2})^{q/(q-1-\nu)}/LD^{2})\rfloor+2,1\}$.
Next, we establish the linear convergence of Algorithm 3. Recall that the pyramidal width is the minimal $\delta>0$ satisfying Lemma B.9 (see [Braun2022-ew, Lemma 2.26] or [Lacoste-Julien2015-gb, Theorem 3] for an in-depth discussion). Under Assumption 4.3, it holds that $\delta\leq D$, and this will be important in establishing convergence rates of Algorithm 3. We make the following assumption:
Assumption 4.3.
The kernel generating distance $\phi$ is $\sigma$-strongly convex.
Theorem 4.4 (Linear convergence of the away-step FW algorithm).
Suppose that Assumptions 3.14.1, and 4.3 hold. Let $\nu>0$ and let $P$ be a polytope and $f$ satisfy the HEB condition with $q>1+\nu$ or $(\nu,q)=(1,2)$. The convergence rate of Algorithm 3 is linear: for all $t\geq 1$
$f(x_{t})-f^{*}\leq\begin{cases}\left(1-\frac{\mu}{32L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ (\nu,q)=(1,2),\\ \frac{1}{(1+\nu)^{\lceil(t-1)/2\rceil}}LD^{2}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)D^{2}c/(\delta/2)^{1+\nu})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\lceil(t-t_{0})/2\rceil\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
where $c:=(q/\mu)^{1/q}$ $D$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and $t_{0}:=\max\{\lfloor\log_{1/(1+\nu)}(L(1+\nu)D^{2}/((\mu/q)^{1/q}\delta/2)^{1+\nu})^{q/(q-1-\nu)}/LD^{2}\rfloor+2,1\}$.
5 Local Linear Convergence for Nonconvex Optimization
In this section, we consider a nonconvex objective function. We establish global sublinear, local sublinear, and local linear convergence (see, for global and local sublinear, Appendices D.1 and D.2). We show that the FW algorithm converges to a minimizer $x^{*}$ when an initial point is close enough to $x^{*}$. We need weak convexity and local quadratic growth for local convergence.
Assumption 5.1.
$f$ is $\rho$-weakly convex and local $\mu$-quadratic growth with $\zeta>0$.
Theorem 5.2 (Local linear convergence of FW algorithm with short step-size or adaptive step-size).
Suppose that Assumptions 3.1 and 5.1 hold. Let $\nu\in(0,1]$ and let $D$ be the diameter. Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$ or the adaptive Bregman step-size strategy (Algorithm 2). Then, if $\rho/\mu<1$, it holds that, for all $t\geq 1$,
where $c=\sqrt{2\mu}/(\mu-\rho)$, $M:=(L(1+\nu))^{1/\nu}$, and $t_{0}:=\max\{\lfloor\log_{\left(1+\nu\rho/\mu\right)/(1+\nu)}((L(1+\nu)(1-\rho/\mu)^{\nu}c^{1+\nu}D^{2}/r^{1+\nu})^{2/(1-\nu)}/LD^{2})\rfloor+2,1\}$.
Finally, we establish the local linear convergence of Algorithm 3. We also assume $\rho\leq L$, which is not restrictive (see also the discussion in Appendix D.2 and Example D.4).
Suppose that Assumptions 3.14.3, and 5.1 hold. Let $\nu\in(0,1]$ and let $P\subset\mathbb{R}^{n}$ be a polytope. The convergence rate of Algorithm 3 with $f$ is linear: if $\rho<\mu\leq L$, for all $t\geq 1$
$f(x_{t})-f^{*}\leq\begin{cases}2\left(1-\frac{\omega}{4L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ \nu=1,\\ 2\left(\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ 1\leq t\leq t_{0},\nu\in(0,1),\\ \frac{(L(1+\nu)(1-\rho/\mu)^{\nu}D^{2}/(\sqrt{\omega}\delta)^{1+\nu})^{2/(1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\left(1-\frac{\rho}{\mu}\right)\lceil(t-t_{0})/2\rceil\right)^{2\nu/(1-\nu)}}=\mathcal{O}(1/t^{2\nu/(1-\nu)})&\text{if}\ t\geq t_{0},\nu\in(0,1),\end{cases}$
where $D$ and $\delta$ are the diameter and the pyramidal width of $P$, respectively, and $\omega:=(\mu-\rho)^{2}/8\mu$, and $t_{0}:=\max\{\lfloor\log_{(1+\nu\rho/\mu)/(1+\nu)}(L^{2}(1+\nu)^{2}D^{4}(1-\rho/\mu)^{2\nu}/(\omega\delta^{2})^{1+\nu})^{1/(1-\nu)}/2LD^{2}\rfloor+2,1\}$.
Without loss of generality, we set $\sigma=1$. If $\rho\leq L$ does not hold, we can use $f(x_{1})-f^{*}\leq(\rho+L)D^{2}$, instead of $f(x_{1})-f^{*}\leq 2LD^{2}$ as the initial bound. Moreover, when $P$ is $\alpha$-strongly convex set and $\phi=\frac{1}{2}\|\cdot\|^{2}$, we also establish local linear convergence. Its proof can be found in Appendix D.5.
6 Numerical Experiments
We compare BregFW and BregAFW, which use the adaptive Bregman step-size strategy (Algorithm 2), with existing algorithms, EucFW, EucAFW, ShortFW, ShortAFW, OpenFW, OpenAFW, and MD (their notation is described in Appendix E). Note that we included OpenAFW only for comparison purposes; there is no established convergence theory for the away-step FW algorithm with the open loop. Indeed, there are no proper drop steps (operations corresponding to line 3 in Algorithm 3), and the favorable properties of the away-step FW algorithm are lost; see [Braun2022-ew]. We use $\beta=0.9$, $\eta=0.9$, $\tau=2$, $\gamma_{\max}=1$, and the termination criterion $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle\leq 10^{-7}$ throughout all experiments.
Figure 1: Log plot of primal and FW gaps on $\ell_{p}$ loss for gas sensor data with $b_{\max}=130$.
Figure 2: Log plot of primal and FW gaps on $\ell_{p}$ loss for gas sensor data with $b_{\max}=200$.
Figure 3: Log plot of primal and FW gaps on phase retrieval for $(m,n)=(1000,10000)$ and $K=2000$.
Figure 4: Log plot of primal and FW gaps on phase retrieval for $(m,n)=(2000,10000)$ and $K=2000$.
6.1 $\ell_{p}$ Loss Problem
We consider the $\ell_{p}$ loss problem [Kyng2015-uo, Maddison2021-nt] to find $x\in P$ such that $Ax\simeq b$, defined by a convex optimization problem $\min_{x\in P}\|Ax-b\|_{p}^{p}$, where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $p>1$. Let $f(x)=\|Ax-b\|_{p}^{p}$. The gradient $\nabla f$ is not Lipschitz continuous on $\mathbb{R}^{n}$ when $p\neq 2$. Furthermore, when $p<2$, $f$ is not $\mathcal{C}^{2}$ but $\mathcal{C}^{1}$. Therefore, for $21<p<$ 2, $\nabla f$ is also not Lipschitz continuous over compact sets. Since $f$ is convex, the pair $(f,\phi)$ is 1-smad with $\phi:=f$.
We consider real-world data using gas sensor data.§§§The dataset is freely available at https://archive.ics.uci.edu/dataset/270/gas+sensor+array+drift+dataset+at+different+concentrations The data size is $(m,n)=(13910,128)$. We use the $\ell_{2}$ constraint $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|^{2}\leq b_{\max}\}$. We compared BregFW with EucFW and OpenFW because ShortFW does not converge. The maximum iteration is 1000 except for OpenFW, and that of OpenFW is 5000. The initial point $x_{0}$ is an all-ones vector. We set $p=1.1$, i.e., $f$ is $\mathcal{C}^{1}$ not $\mathcal{C}^{2}$. Figures 4 and 4 show the primal and FW gaps per iteration and gaps per second for $b_{\max}=130$ and $b_{\max}=200$, respectively. Because $\nabla f$ is not Lipschitz continuous on $P$ from lacking $\mathcal{C}^{2}$, ShortFW and EucFW do not converge. The primal gap and the FW gap by BregFW are the smallest among these step-size strategies because only BregFW has a theoretical guarantee for this problem.
6.2 Phase Retrieval
We are interested in phase retrieval, which involves finding a signal $x\in\mathbb{R}^{n}$ such that $|\langle a_{i},x\rangle|^{2}\simeq b_{i}$, $i=1,\ldots,m$, where $a_{i}\in\mathbb{R}^{n}$ describes the model and $b\in\mathbb{R}^{m}$ is a vector of measurements. Phase retrieval arises in many fields of science and engineering, such as image processing [Candes2015-mk], X-ray crystallography [Patterson1934-sg, Patterson1944-nm], and optics [Shechtman2015-ic]. To achieve the goal of phase retrieval, we focus on the nonconvex optimization problem $\min_{x\in P}f(x):=\frac{1}{4}\sum_{i=1}^{m}\left(|\langle a_{i},x\rangle|^{2}-b_{i}\right)^{2}$, where $P\subset\mathbb{R}^{n}$ is a compact convex set. Let $\phi$ be defined by $\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2}$, which is $\sigma$-strongly convex for $\sigma\leq 1$. For any $L$ satisfying $L\geq\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right)$, the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Bolte2018-zt, Lemma 5.1]. In addition, $\nabla^{2}f(x)+\rho I$ is positive semidefinite for $\rho\geq\sum_{i=1}^{m}\|a_{i}\|^{2}|b_{i}|$, i.e., $f$ is $\rho$-weakly convex. Therefore, $L\geq\rho/\sigma$ holds with $\sigma=1$.
We use a $K$-sparse polytope as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{1}\leq K,\|x\|_{\infty}\leq 1\}$. We compared BregFW with EucFW, ShortFW, and OpenFW. The maximum iteration is 1000 except for OpenFW, and that of OpenFW is 5000. We generated $a_{i}$, $i=1,\ldots,m$ from an i.i.d. normal distribution and normalized it to have norm 1 (random seed 1234). We also generated $x^{*}$ from an i.i.d. uniform distribution in $[0,1]$ and normalized $x^{*}$ to have sum 1. The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$. Figures 4 and 4 show the primal and FW gaps per iteration and gaps per second for $(m,n)=(1000,10000)$ and $K=2000$ and for $(m,n)=(2000,10000)$ and $K=2000$, respectively. The primal gap and the FW gap by the adaptive Bregman step-size strategy are the smallest among these step-size strategies. In both cases, BregFW stopped before the 1000th iteration.
7 Conclusion and Future Work
Summary:
We propose FW algorithms with an adaptive Bregman step-size strategy for a smooth adaptable function, which does not require Lipschitz continuity of its gradient. We have established sublinear to linear convergence rates depending on various assumptions for convex and nonconvex optimization. We also propose a variant of the away-step FW algorithm using Bregman distances over polytopes. For the away-step FW algorithm, we have also established global convergence for convex optimization under the HEB condition and its local linear convergence for nonconvex optimization under the local quadratic growth condition. Numerical experiments show that our algorithms outperform existing algorithms.
Limitation and future work:
Our step-size strategy for a smooth adaptable function requires the value of $\nu$, which can be estimated, because we use the Bregman distance. Theoretically, linear convergence of the FW algorithm always holds for convex optimization if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$. For $q>1+\nu$, if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$, linear convergence holds; otherwise, sublinear convergence. The condition $q=1+\nu$, which is derived from the growth of $f$ and $\phi$, is natural because $\phi$ has a similar structure to $f$. On the other hand, for nonconvex optimization, we assume the local quadratic growth condition, which is the special case of the HEB condition with $q=2$. Since this is derived from the weak convexity of $f$, relaxing the quadratic growth condition may not be possible. In that regard, one direction is to consider an extension using the difference of convex functions (DC) optimization. Recently, Maskan et al. [Maskan2025-oa] incorporated DC optimization into FW algorithms. While they established sublinear convergence, our algorithms may exploit DC optimization techniques.
Acknowledgement
Appendix A Results from Smooth Adaptable Property
A.1 Properties of Bregman Distances